MTA Monitoring Report 9/15/00 - 9/21/00

Archive of all weekly reports

Archive of Monthly reports

************ Radiation *************

Event Sunday 9/16 cleared two hour threshold. Breif event, no action taken.
Air Force Kp violations Sunday 9/17 (7.33) and Tuesday 9/19 (6.0). Total of 9 hours of violation time. No action taken.
Costello Kp Violation 9/16 (6.66) and 9/17 (7.33) 15 hours Kp > 5.5. No action taken.

************ CTI trend *************

Average of ACIS-I CCDs (MnKa) slope: 2.481e-8 (CTI/day) Detrended slope 1.374e-8 (CTI/day)

*********** ACIS Focal Temperature ********

For this period, there were 2 peaks.
       Day (DOY)       Temp (C)        Width (DAY)
        260.3           -111.5          0.48
        262.9           -111.3          0.37

************** MTA Problem Tracking *****************

No new or reopened problem was reported in this period. Following msids are still coming (All eclipse related except AWD3TQI)

************** SIM Movement ************

7 TSC moves this period average time/step 0.00139 s nominal Max Pulse Width Modulation for each move see SIM monitoring

************* Telemetry ***********

Violation related to OTG moves and eclipse operations.
9/15, battery msids (EB1DIC, EOEB1CIC, EOEB2CIC, EOEB3CIC) showed a large oscillation into red violations.
9/18, many of oba msids (esp. OBA cone/OBA fwd bulkhead) registered a sharp peak (and a red or yellow violation).
Spacecraft electronics, OOBAPWR (OBA COMPUTED TOTAL POWER) also showed a huge spike to 800 level (red upper limits: 300.0, usually around 0-200) ELBI (LOAD BUS CURRENT) showed a similar spike.
9/20, Grating voltages 4MP5AV,4MP5BV,4MP28AV, and 4MP28BV displayed a large spike. Nominally, they are all zeros (current settings), but jumped to the other setting limits (5V, 5V, 28V, 28V, respectively).

************ Trending *************

Spacecraft Electronics A, CUSOA28V (+28V USO A) jumped from steady 23.85 to steady 24.5 for the last month or so. We need to see if this continnues now that eclipses are done.

*********** Recent Observations ************

The ACA Problem is in the specified handling of guide star positional errors. This issue in not signifcant to the final product and so is being worked at low priority.


10      ACIS-S  HETG    AR LAC-5 FIRST QUAD     ok	    OK
774     ACIS-S  HETG    0509-67.5               ok	    OK
11      ACIS-S  HETG    AR LAC-6 LAST QUAD      ok	    OK
2305    ACIS-S  HETG    TAU SCO                 ok          PROBLEM
638     ACIS-S  HETG    TAU SCO                 ok          OK
648     ACIS-S  HETG    SS CYG                  ok          PROBLEM
9       ACIS-S  HETG    AR LAC-4 ECLIPSE        ok          OK
104     ACIS-S  HETG    4U 1626-67              ok          OK
7       AICS-S  HETG    AR LAC-2 FIRST QUAD     ok          OK
8       ACIS-S  HETG    AR LAC-3 LAST QUAD      ok          PROBLEM